# VisTR: Visualizations as Representations for Time-series Table Reasoning

### v2

@@ Abstract
Table reasoning involves transforming natural language questions into corresponding answers based on the provided data table. Recent research exploits large language models (LLMs) to facilitate table reasoning, which however struggle with pattern recognition and lack support for visual-based pattern exploration.To address these limitations, we propose VisTR, a framework that leverages visualizations as representations to facilitate data pattern recognition and support cross-modal exploration. We describe VisTR as a process consisting of four major modules: 1) visualization alignment that utilizes multimodal LLMs to align visualizations across various modalities, including chart, text, and sketch; 2) visualization referencing that decomposes a table into multifaceted visualization references that comprehensively represent the table; 3) visualization pruning that incorporates data and retrieval pruning to excise visualization references with poor information and enhance retrieval efficiency; and 4) visualization interaction that offers an interactive visual interface with multimodal interactions for user-friendly table reasoning. Quantitative evaluation with existing multimodal LLMs demonstrates the effectiveness of the alignment model in cross-modal visualization pairings. We further illustrate the applicability of the proposed framework in various time-series table reasoning and exploration tasks.
keywords:
Table reasoning, visualization representation, multimodal LLMs
[…]
@@
VisTR leverages visualizations as the representations between a table and user intentions. The proposed framework advances existing LLM-based table reasoning methods by supporting data change pattern recognition (Q1), improving pattern recognition for long tables (Q2), and enabling visual-based pattern exploration (Q3).
1 Introduction
@@
Table reasoning is a crucial natural language processing task that involves automatically extracting and inferring information from tables to answer specific questions[12, 70]. This task supports a wide range of applications such as fact verification[6] and question answering[26], where each question-table pair constitutes an instance for reasoning[69]. Given the superior commonsense and logical reasoning capabilities of large language models (LLMs)[63], utilizing them for table reasoning has emerged as the predominant approach[5, 12]. Existing methods can be broadly categorized into generic table reasoning, where pre-trained models or LLMs directly generate answers, often enhanced by techniques such as designing prompts[5] to guide iterative reasoning or employing flexible tabular formats[57, 69]; and program-aided table reasoning that utilizes executable query languages (e.g., SQL, SPARQL, or Python) as the bridge between user queries and answers derived from the table[65, 9, 19].
Despite their advancements, there exist several limitations. First, these methods struggle with pattern recognition, which involves the automatic identification and extraction of meaningful patterns within data, such as trends in time-series data[35, 8]. For instance, Q1 in Figure VisTR: Visualizations as Representations for Time-series Table Reasoning illustrates a scenario where users seek to examine the price trend of Apple during March. Existing LLM-based methods cannot correctly capture and respond to the data change patterns within the corresponding time period. The challenge becomes more pronounced when dealing with large-size tables, as LLMs are known to encounter ‘drift’ issues, where models lose information in the middle of long contexts[4]. This loss of context can result in errors during the reasoning process, making it difficult to extract patterns accurately, as illustrated in Figure VisTR: Visualizations as Representations for Time-series Table Reasoning(Q2). Last, the methods lack support for visual-based pattern exploration, making it challenging for users to iteratively engage with data tables, particulalry for lay users. Rather than expressing their intentions precisely in natural language, laymen would prefer to use multimodal inputs like hand-drawn sketches or charts[41]. An example is shown in Figure VisTR: Visualizations as Representations for Time-series Table Reasoning(Q3), where a user may not know the term “two-valley” but can easily recognize this visual pattern from the collected bar chart. However, existing LLM-based methods do not support such visual-based exploration, which are increasingly in demand.
To overcome the limitations of existing LLM-based table reasoning methods, we introduce a novel framework that incorporates visualization as representation into multimodal LLMs (MLLMs). Our key idea is to transform data tables into a set of meaningful and insightful visualizations, termed visualization references, which serve as a concise data representation to capture data insights and align with human intentions. This framework offers several advantages that address the shortcomings of existing LLM-based table reasoning methods, effectively enhancing the table exploration process. First, visualization references can help recognize the short- and long-term data change patterns. For instance, when recognizing price trend in the given data table (Figure VisTR: Visualizations as Representations for Time-series Table Reasoning(Q1)), the corresponding visualization reference could capture the “two-peak” price trend of Apple during March. Second, regardless of the length of data facets, the visualization references can capture and retain the data change patterns in fixed-size images. Adopting fixed-size visualization references as the representations of data facets can effectively circumvent the ‘drift’ issues encountered during reasoning. For Q2 in Figure VisTR: Visualizations as Representations for Time-series Table Reasoning, the stock price change over a year can be represented as a single visualization, where the price bottom can be easily located in this reference. Moreover, leveraging visualizations as a representation offers new modalities for interacting with the tabular data, aligning more intuitively with human intentions[32]. Combined with multimodal input, including text and chart (as shown in Figure VisTR: Visualizations as Representations for Time-series Table Reasoning(Q3)), it becomes easier to explore data tables than directly observing the raw table.
We illustrate the effectiveness of VisTR on time-series data tables, which often feature large table sizes and complex data change patterns. In this context, the proposed framework also presents several challenges to be addressed. First, time-series data often contain a variety of long- and short-term patterns, which should be comprehensively captured by the visualization references. To ensure the visualization references are comprehensive enough to accurately represent a data table, thereby meeting the diverse analytical needs, poses the first challenge (C1). Second, to effectively align textual descriptions, visualization images, and hand-drawn sketches to bridge user intentions with visualization references, meanwhile enabling multimodal interactions presents a challenge (C2). Existing models for vision-language tasks such as chart question answering (CQA)[24, 68] are not specifically designed to meet this requirement, thus being less effective in identifying data change patterns. Last, to comprehensively represent a table, a large number of visualization references will be generated. Effectively indexing the visualization references and supporting efficient visualization retrieval for timely user response present a significant challenge (C3).
We design a prototype system that encompasses four main modules aimed at addressing the challenges C1--C3. First, the visualization referencing module decomposes an input table into data facets and processes the facets into corresponding visualization references that comprehensively capture the essence of the table. To ensure a comprehensive representation of data patterns in tabular data, we adopt a multifaceted process to get meaningful facets and then transform each facet into a fixed-size visualization reference (addressing C1). Second, in the visualization alignment module, we fine-tune a MLLM to connect visualizations of various modalities including charts, text, and sketches (addressing C2). Specifically, we construct a new cross-modal visualization dataset using data augmentation for chart-text pairing and user labeling for chart-sketch pairing, taking into account user intentions. We further leverage the CLIP model[48] enriched with two-level cross-entropy loss and bidirectional hinge-based triplet ranking loss to enhance the alignment. Next, the visualization pruning module leverages visualization pruning to filter out less informative visualization references, and visualization indexing to enhance the efficiency of the retrieval process for the embeddings of visualization references stored in a vector database (addressing C3). Finally, we design a visualization interface that enables users’ exploration of time-series data tables through multimodal interactions. Here, we incorporate a “decomposing-execution-filling” strategy to integrate the output generated by VisTR with the input table. Specifically, the user input is initially decomposed into visualization retrieval tasks, which are executed in the vector database, and their corresponding text descriptions are extracted by the fine-tuned MLLM. Subsequently, these descriptions are filled back into the input table to generate textual responses for the user. Quantitative evaluations demonstrate the effectiveness of the fine-tuned MLLM in aligning different modalities. Additionally, case studies illustrate how the proposed framework addresses the limitations of existing LLM-based table reasoning methods through various time-series table reasoning and visual-based exploration tasks. Dataset, fine-tuned model, and evaluation results are released at https://osf.io/yun9d/.
@@
(1) Framework. We introduce a framework that leverages visualizations as the representations to enhance table reasoning, to facilitate data change pattern recognition and visual-based exploration.
(2) Multimodal Visualization Alignment. We design an effective MLLM that learns a joint embedding space across three visualization modalities - charts, text, and hand-drawn sketches. The model aligns multimodal user intentions with visualization references, allowing users to explore and analyze tables through various interactions.
(3) VisTR System. We develop a prototype system that supports the exploration and reasoning of large-scale and multiple time-series tables through intuitive user interactions. Various case studies and quantitative results prove the effectiveness and applicability of VisTR.
2 Related Work
Table Reasoning. Table reasoning integrated with natural language (NL) interfaces enables lay users to explore tabular data effortlessly, supporting complex tasks such as fact verification[6] and question answering[26]. Existing LLM-based table reasoning methods can be categorized into generic table reasoning and program-aided table reasoning[57]. Generic reasoning methods employ LLMs to interpret and respond to free-form NL queries by generating answers directly from tables[5]. However, due to the complexity of table structure, these methods struggle with large tables and offer limited explainability because of their end-to-end nature. Program-aided table reasoning utilizes executable query languages, such as SQL, to form a structured and accountable bridge between user quires and the derived answers[14, 37, 25]. Recent research has mostly focused on leveraging LLMs to transform complex user utterances into SQL queries with data context awareness[65], using prompt engineering[30], chain-of-thought[57, 66], and pre-training techniques[56]. However, program-aided approaches encounter intricate limitations due to their reliance on query language. First, SQLs primarily cater to numerical queries from tables but lack support of analytical functionality[29], overlooking data change patterns. Second, when confronted with large tables, LLMs may encounter drift issues[4] while accessing information in the middle of long contexts. Moreover, in many scenarios, relying solely on textual inputs limits the expression of non-verbal intentions, such as hand-drawn sketches, reducing the flexibility of user interactions.
@@ 2 Related Work
Multimodal Visualization Alignment. Given that visualizations are inherently stored as images, we opt to leverage emerging MLLMs to embed visualization images and facilitate multimodal user interactions. In recent years, significant advancements have been made in utilizing vision-language models for chart-based tasks, such as CQA[42, 43, 24], chart retrieval[60] and generation[50, 61]. Though CQA models can provide answers for user utterances, these models only support text-chart interactions, lacking the capability for cross-modal interactions involving multiple modalities such as text, chart, and sketch preferred by laymen. Furthermore, these models are not specifically designed for effectively recognizing data change patterns (see Sect. 6.1), a gap that this work seeks to fill in the context of table reasoning.
@@
To this end, we choose to construct a novel vision-language model with an innovative visualization alignment strategy. Most existing multimodal data alignment methods rely on contrastive language-image pre-training (CLIP)[48], which has demonstrated its adaptability and proficiency across various tasks[71, 34, 72], highlighting the effectiveness of joint embedding spaces. However, training an effective CLIP-based alignment model for visualizations is non-trivial. First, this task involves aligning more than two modalities, including text, hand-drawn sketches, and various types of charts, whereas the original CLIP model was designed to support text-image alignment. To encompass more modalities, we adopt joint embedding strategies that leverage charts as naturally aligned supervision and emergent alignments to unify other modalities[20]. This approach allows us to prepare paired data only for chart-text and chart-sketch, rather than other pairings such as sketch-text. Second, the effectiveness of CLIP is largely attributed to its extensive training on vast (natural) image-text datasets, whereas there is a lack of suitable training datasets for the task in this work. Efforts have been made to construct datasets for promoting deep learning models for chart-text (e.g.,[28, 42]), chart-chart (e.g.,[40, 33, 64]), chart-sketch (e.g.,[41, 52, 38]) similarity measurement. However, these different modalities are not inherently unified in one model, and many of these datasets are not specifically designed for recognizing data change patterns. To address this gap, we utilize data augmentation and user labeling strategies to supplement existing datasets in training the multimodal visualization alignment model.
@@
Visualization Storage, Indexing, and Retrieval. To comprehensively represent tabular data, our approach will generate a large number of visualization references and encode them into vector representations, thus raising new challenges for efficiently storing, indexing, and retrieving them. With the recent advancements in Vector Database Management Systems (VDBMS)[46], there is potential to meet our demand. VDBMSs are specifically designed for managing high-dimensional data, such as embeddings, offering sophisticated storage, indexing, query processing, and query optimization capabilities[3]. Their ability to perform efficient similarity searches in high-dimensional spaces makes them particularly suited for tasks involving the retrieval of embeddings based on content or semantic similarity. This supports various applications such as information retrieval, e-commerce, and recommendation platforms[22, 55]. Existing VDBMS can be roughly categorized into two types: native systems and extended systems[46]. The former, such as Chroma[1] and EuclidesDB[2], are tailored for vector data management, while the latter, such as AnalyticDB-V[58], are built on top of an existing data management system.
@@
3 The Visualization Enhanced Framework
@@
The visualization enhanced framework consists of four modules: visualization alignment, visualization referencing, visualization pruning, and visualization interaction, as shown in Figure 1.
@@
Visualization Alignment. As the cornerstone of our framework, a MLLM is fine-tuned to establish a single joint embedding space for various modalities, including chart ($\mathbb{C}$), text ($\mathbb{T}$), and sketch ($\mathbb{S}$). To address the issue of a lack of suitable training data, we first curate a new dataset with chart-text pairings augmented from existing Chart-to-Text dataset[28], and chart-sketch pairings through user labeling. Utilizing the pre-trained CLIP[48], we fine-tune a multimodal LLM on our composite dataset, employing a Transformer[49] for text modality, and ViT[13] encoders for chart and sketch modalities, respectively. To achieve alignment across modalities within the joint embedding space, we implement a bidirectional hinge-based triplet ranking loss function and a two-level cross-entropy loss function. Through this process, we successfully align the three modalities and yield strong performance across four evaluation metrics surpassing a threshold of 0.85. A detailed description of the module is provided in Sect. 4.
@@
Visualization Referencing. In the visualization referencing module, we generate visualization references that comprehensively capture key data patterns inherent in an input table. This module employs a two-stage approach that first decomposes the input table $D$ into data facets, which are then mapped into fixed-size visualization references. The description of this module is detailed in Sect. 5.1.
@@
Visualization Pruning. In the visualization pruning module, we introduce pruning and indexing techniques to enable timely visualization reference retrieval. Specifically, visualization pruning works by filtering out less informative visualization references by measuring the similarity between visualization vectors using Euclidean distance and applying a pruning threshold. This process reduces the number of stored references while preserving essential data change patterns. Concurrently, visualization indexing leverages an open-source vector database, Chroma[1], to efficiently store and retrieve visualization references. Here, we use the fine-tuned MLLM to transform all the pruned visualization references into 512-dimensional vectors, which are then indexed by Chroma for rapid and accurate follow-up retrieval. This combination of pruning and indexing ensures VisTR can swiftly manage and retrieve massive visualization references with precision. This module is detailed in Sect. 5.2.
@@
Visualization Interaction. In the visualization interaction module, we introduce a “decomposing-execution-filling” strategy to enable the connection of multimodal user queries with the stored embeddings of visualization references. Specifically, the user query is first decomposed into a visualization retrieval task. Second, the retrieval task is executed in Chrome to get a retrieved visualization reference. Afterward, with the help of the fine-tuned MLLM, the corresponding trend word of the retrieved visualization reference is generated and then filled into the textual response to the user. An interactive visual interface with multiple views and rich interactions is also developed to support the iterative table reasoning process, transitioning from abstract to detailed exploration, as detailed in Sect. 5.3.
4 Multimodal Visualization Alignment

### v3

# mO Yuyu[#1]

@@ Abstract
Time-series table reasoning interprets temporal patterns and relationships in data to answer user queries. Despite recent advancements leveraging large language models (LLMs), existing methods often struggle with pattern recognition, context drift in long time-series data, and the lack of visual-based reasoning capabilities. To address these challenges, we propose VisTR, a framework that places visualizations at the core of the reasoning process. Specifically, VisTR leverages visualizations as representations to bridge raw time-series data and human cognitive processes. By transforming tables into fixed-size visualization references, it captures key trends, anomalies, and temporal relationships, facilitating intuitive and interpretable reasoning. These visualizations are aligned with user input, i.e., charts, text, and sketches, through a fine-tuned multimodal LLM, ensuring robust cross-modal alignment. To handle large-scale data, VisTR integrates pruning and indexing mechanisms for scalable and efficient retrieval. Finally, an interactive visualization interface supports seamless multimodal exploration, enabling users to interact with data through both textual and visual modalities. Quantitative evaluations demonstrate the effectiveness of VisTR in aligning multimodal inputs and improving reasoning accuracy. Case studies further illustrate its applicability to various time-series reasoning and exploration tasks.
Index Terms:
Table reasoning, visualization representation, multimodal LLMs.
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
@@
Figure 1: VisTR leverages visualizations as the representations to bridge the gap between time-series data and user queries. VisTR enhances existing LLM-based table reasoning methods by enabling robust data change pattern recognition (Q1), improving pattern recognition for long-term series (Q2), and facilitating visual-based pattern exploration (Q3).
Time-series tabular data are ubiquitous, forming the backbone of decision-making processes across domains such as finance, healthcare, and logistics[1, 2]. These datasets record sequential information over time, enabling the discovery of temporal trends and interesting relationships that drive actionable insights[3, 4]. Time-series table reasoning, a key task in this context, focuses on interpreting these patterns to answer user queries, such as “What is the price trend of Apple during March?” (see Fig. 1-Q1). Intuitively, effectively addressing this task requires a comprehensive understanding of not only the raw tabular data but also the temporal trends, interrelationships among data points, and alignment with the specific context of user queries [5].
@@
Recently, large language models (LLMs) have demonstrated promising capabilities in extending their applications to table reasoning tasks [6, 7, 8]. These methods can be broadly categorized into: (1) generic table reasoning, where pre-trained models or LLMs directly generate answers, often enhanced by techniques such as prompt design [9, 10], iterative reasoning, or flexible tabular formats [11, 12]; and (2) program-aided table reasoning which employs executable query languages (e.g., SQL, SPARQL, or Python) to bridge user queries with answers derived from tabular data [13, 14, 15].
Limitations of Existing Solutions. Despite these advances, existing solutions exhibit limitations, as illustrated in Fig. 1.
(1) Limited Pattern Recognition. LLMs often struggle to identify and interpret temporal data patterns, such as trends, peaks, or valleys [1, 16]. For example, in Q1 of Fig. 1, an LLM-based approach misinterprets Apple’s stock prices as exhibiting a “declining trend” rather than recognizing the correct two-peak trend.
(2) Challenges with Long Time-series Data. LLMs often lose focus and context when reasoning over long time-series data, leading to errors caused by “LLM drift” [17]. In Q2 of Fig. 1, an LLM-based method erroneously identifies Google’s stock price bottom as occurring “at the beginning” rather than pinpointing the correct date.
(3) Inability to Perform Visual-based Reasoning. Existing LLMs cannot effectively reason over external visual references, limiting their ability to align time-series data with visual patterns [18]. For example, in Q3 of Fig. 1, an LLM-based method fails to determine whether Tesla’s stock price exhibits a similar pattern to a user-provided chart, demonstrating the gap in visual-based reasoning capabilities.
Rethinking: How Can Visualization Help? Visualizations, such as line and bar charts, offer an intuitive means of representing time-series data by highlighting trends, outliers, and temporal relationships [19, 20]. They act as a cognitive bridge, enabling humans to discern patterns that might otherwise remain obscured in raw time-series data.
Despite their potential, existing table reasoning methods largely overlook the use of visualizations as proxies for reasoning. Instead, they rely heavily on text-based reasoning, which struggles to emulate human-like cognitive processes and often fails in tasks requiring complex pattern recognition and context alignment with user intentions.
Our Proposal. Inspired by the ability of visualizations to simplify complex patterns and align with human cognitive processes, we propose a new framework that integrates visualizations as representations into multimodal large language models (MLLMs). Our goal is to achieve the best of both worlds by combining the interpretability and intuitive pattern recognition offered by visualizations with the reasoning and generative capabilities of LLMs, thereby enhancing both the accuracy and usability of time-series table reasoning systems. The key idea of our framework is to transform time-series data into a set of meaningful and insightful visualizations, termed visualization references. These references act as concise representations of the data, capturing essential insights such as trends, anomalies, and temporal relationships, while aligning closely with user intentions for reasoning tasks.
Framework Overview. At its core, the framework transforms input tables into fixed-size visualization references that capture key trends, anomalies, and temporal relationships, enabling a comprehensive and interpretable representation of both long- and short-term data patterns. To achieve this, the framework begins by decomposing tables into smaller, meaningful data facets, each representing a subset of the table’s temporal or structural characteristics. These facets are then converted into visualizations, ensuring a rich and diverse representation of the data. Second, to align user intentions with these visualization references, the framework incorporates a fine-tuned MLLM that constructs a unified embedding space for multimodal inputs, including charts, text, and hand-drawn sketches. A novel cross-modal dataset is developed to enhance alignment precision, combining augmented chart-text pairs with user-labeled chart-sketch pairs. By extending the CLIP model [21] with advanced loss functions, e.g., a two-level cross-entropy loss and a bidirectional triplet ranking loss, the framework ensures accurate and consistent alignment across modalities, bridging the gap between user inputs and relevant visualizations. Third, given the large-scale visualization references generated from extensive datasets, the framework incorporates a pruning and indexing mechanism to enhance scalability and retrieval efficiency. Less informative visualizations are filtered out, while the remaining embeddings are indexed in a vector database, enabling fast and accurate retrieval during user interactions. Finally, the framework integrates these components into an intuitive visualization interface supporting seamless multimodal interaction. Using a “decomposing-execution-filling” strategy, user queries are broken into retrieval tasks, linked to textual descriptions generated by the fine-tuned MLLM, and reintegrated into the table for final responses. This interface enables interactive exploration through charts, text, and sketches. Our framework offers several key advantages over existing LLM-based table reasoning methods.
(1) Enhanced Pattern Recognition Across Time Scales. Visualization references effectively capture both short- and long-term data change patterns, making it easier to identify nuanced temporal trends. For example, in Fig. 1(Q1), the “two-peak” price trend of Apple during March is clearly represented, overcoming the limitations of purely text-based approaches.
(2) Mitigating Context Drift in Long Time-series. Fixed-size visualization references condense data patterns, mitigating context drift issues in long time-series reasoning. For example, in Fig. 1(Q2), a year-long stock price change is condensed into a single visualization, making it easy to pinpoint the price bottom without losing contextual accuracy.
(3) Intuitive Multimodal Interaction. Integrating visualizations as representations enables natural, multimodal data exploration by aligning with human cognitive processes. For example, Fig. 1(Q3) shows how combining textual descriptions with charts improves query interpretation and reduces reliance on raw time-series data [22].
@@
  • Framework. We propose VisTR that leverages visualizations as representations to enhance time-series table reasoning. This new framework enables robust pattern recognition and intuitive exploration for time-series data.
  • Multimodal Visualization Alignment. We finetune an effective MLLM that learns a joint embedding space across three modalities, i.e., charts, text, and hand-drawn sketches. This alignment bridges user intentions with visualization references, enabling users to explore time-series tables through multimodal interactions.
  • Prototype Construction. We implement a prototype system that facilitates reasoning and exploration of large-scale time-series tables.
  • Evaluation. We conduct extensive evaluations, including quantitative experiments and case studies, to demonstrate the effectiveness and usability of VisTR. All datasets, fine-tuned models, and evaluation results are available at https://osf.io/yun9d/.
II Related Work
II-A Table Reasoning
Table reasoning integrated with natural language (NL) interfaces enables lay users to explore tabular data effortlessly, supporting complex tasks such as fact verification[23] and question answering[24]. Existing LLM-based table reasoning methods can be categorized into generic table reasoning and program-aided table reasoning[11]. Generic reasoning methods employ LLMs to interpret and respond to free-form NL queries by generating answers directly from tables[9]. However, due to the complexity of table structure, these methods struggle with large tables and offer limited explainability because of their end-to-end nature. Program-aided table reasoning utilizes executable query languages, such as SQL, to form a structured and accountable bridge between user quires and the derived answers[25, 26, 27]. Recent research has mostly focused on leveraging LLMs to transform complex user utterances into SQL queries with data context awareness[13], using prompt engineering[28], chain-of-thought[11, 29], and pre-training techniques[30]. However, program-aided approaches encounter intricate limitations due to their reliance on query language. First, SQLs primarily cater to numerical queries from tables but lack support of analytical functionality[16], overlooking data change patterns. Second, when confronted with large tables, LLMs may encounter drift issues[17] while accessing information in the middle of long contexts. Moreover, in many scenarios, relying solely on textual inputs limits the expression of non-verbal intentions, such as hand-drawn sketches, reducing the flexibility of user interactions.
@@ II-A Table Reasoning
II-B Multimodal Visualization Alignment
@@
Given that visualizations are inherently stored as images, we opt to leverage emerging MLLMs to embed visualization images and facilitate multimodal user interactions. In recent years, significant advancements have been made in utilizing vision-language models for chart-based tasks, such as CQA [35, 36, 37], chart retrieval [38] and generation [39, 40]. Though CQA models can provide answers for user utterances, these models only support text-chart interactions, lacking the capability for cross-modal interactions involving multiple modalities such as text, chart, and sketch preferred by laymen [41]. Furthermore, these models are not specifically designed for effectively recognizing data change patterns (see Sect. VI-A), a gap that this work seeks to fill in the context of table reasoning. To this end, we choose to construct a novel vision-language model with an innovative visualization alignment strategy. Most existing multimodal data alignment methods rely on contrastive language-image pre-training (CLIP) [21], which has demonstrated its adaptability and proficiency across various tasks[42, 43, 44], highlighting the effectiveness of joint embedding spaces. However, training an effective CLIP-based alignment model for visualizations is non-trivial. First, this task involves aligning more than two modalities, including text, hand-drawn sketches, and various types of charts, whereas the original CLIP model was designed to support text-image alignment. To encompass more modalities, we adopt joint embedding strategies that leverage charts as naturally aligned supervision and emergent alignments to unify other modalities[45]. This approach allows us to prepare paired data only for chart-text and chart-sketch, rather than other pairings such as sketch-text. Second, the effectiveness of CLIP is largely attributed to its extensive training on vast (natural) image-text datasets, whereas there is a lack of suitable training datasets for the task in this work. Efforts have been made to construct datasets for promoting deep learning models for chart-text (e.g.,[46, 35]), chart-chart (e.g.,[47, 48, 49]), chart-sketch (e.g.,[41, 50, 51]) similarity measurement. However, these different modalities are not inherently unified in one model, and many of these datasets are not specifically designed for recognizing data change patterns. To address this gap, we utilize data augmentation and user labeling strategies to supplement existing datasets in training the multimodal visualization alignment model.
II-C Visualization Storage, Indexing, and Retrieval
@@
To comprehensively represent tabular data, our approach will generate a large number of visualization references and encode them into vector representations, thus raising new challenges for efficiently storing, indexing, and retrieving them. With the recent advancements in Vector Database Management Systems (VDBMS)[52], there is potential to meet our demand. VDBMSs are specifically designed for managing high-dimensional data, such as embeddings, offering sophisticated storage, indexing, query processing, and query optimization capabilities[53]. Their ability to perform efficient similarity searches in high-dimensional spaces makes them particularly suited for tasks involving the retrieval of embeddings based on content or semantic similarity. This supports various applications such as information retrieval, e-commerce, and recommendation platforms[54, 55]. Existing VDBMS can be roughly categorized into two types: native systems and extended systems[52]. The former, such as Chroma[56] and EuclidesDB[57], are tailored for vector data management, while the latter, such as AnalyticDB-V[58], are built on top of an existing data management system.
@@
III The Visualization Enhanced Framework
@@
The visualization enhanced framework consists of four modules: visualization alignment, visualization referencing, visualization pruning, and visualization interaction, as shown in Fig. 2.
@@
Visualization Alignment. As the cornerstone of our framework, a MLLM is fine-tuned to establish a single joint embedding space for various modalities, including chart ($\mathbb{C}$), text ($\mathbb{T}$), and sketch ($\mathbb{S}$). To address the issue of a lack of suitable training data, we first curate a new dataset with chart-text pairings augmented from existing Chart-to-Text dataset[46], and chart-sketch pairings through user labeling. Utilizing the pre-trained CLIP[21], we fine-tune a multimodal LLM on our composite dataset, employing a Transformer[59] for text modality, and ViT[60] encoders for chart and sketch modalities, respectively. To achieve alignment across modalities within the joint embedding space, we implement a bidirectional hinge-based triplet ranking loss function and a two-level cross-entropy loss function. Through this process, we successfully align the three modalities and yield strong performance across four evaluation metrics surpassing a threshold of 0.85. A detailed description of the module is provided in Sect. IV.
@@
Visualization Referencing. In the visualization referencing module, we generate visualization references that comprehensively capture key data patterns inherent in an input table. This module employs a two-stage approach that first decomposes the input table $D$ into data facets, which are then mapped into fixed-size visualization references. The description of this module is detailed in Sect. V-A.
@@
Visualization Pruning. In the visualization pruning module, we introduce pruning and indexing techniques to enable timely visualization reference retrieval. Specifically, visualization pruning works by filtering out less informative visualization references by measuring the similarity between visualization vectors using Euclidean distance and applying a pruning threshold. This process reduces the number of stored references while preserving essential data change patterns. Concurrently, visualization indexing leverages an open-source vector database, Chroma[56], to efficiently store and retrieve visualization references. Here, we use the fine-tuned MLLM to transform all the pruned visualization references into 512-dimensional vectors, which are then indexed by Chroma for rapid and accurate follow-up retrieval. This combination of pruning and indexing ensures VisTR can swiftly manage and retrieve massive visualization references with precision. This module is detailed in Sect. V-B.
@@
Visualization Interaction. For visualization interaction, we introduce a “decomposing-execution-filling” strategy to enable the connection of multimodal user queries with the stored embeddings of visualization references. Specifically, the user query is first decomposed into a visualization retrieval task. Second, the retrieval task is executed in Chroma to get a retrieved visualization reference. Afterward, with the help of the fine-tuned MLLM, the corresponding trend word of the retrieved visualization reference is generated and then filled into the textual response to the user. An interactive visual interface with multiple views and rich interactions is also developed to support the iterative table reasoning process, transitioning from abstract to detailed exploration, as detailed in Sect. V-C.
IV Multimodal Visualization Alignment